# Radiative lifetime of the state in RaF with relevance to laser cooling

### v1

@@ IV Conclusions
In this work, the measurement of the radiative lifetime of the () state in RaF is presented. The lifetime was measured via delayed multi-step ionization using the CRIS experiment at the CERN-ISOLDE radioactive ion beam facility. The lifetime is determined to be ns. Using this value, the radiative decay rate s is extracted, which further determines the natural linewidth of the transition MHz and the maximum photon scattering rate s. The validity of the lifetime measurement in RaF is benchmarked via a measurement of the lifetime of the state in Fr using the same technique. The extracted lifetime for the state in Fr at 83(3) ns in this work is in agreement with the literature value.
The quantities reported in this work are of direct relevance to simulations for efficient laser cooling of RaF. Laser-cooled and trapped neutral molecules are envisioned as sensitives probes for future searches of new physics, and thanks to its exceptionally diagonal Franck-Condon matrix, RaF is expected to scatter 10 photons using only 2 cooling lasers, while using 3 cooling lasers allows scattering an order of magnitude more photons than other laser-cooled molecules. While the photon-scattering acceleration is lower in RaF compared to the lighter homoelectronic molecules, it is shown that the difference becomes insignificant for a molecular source that produces a sufficiently slow beam (50 m/s).
@@ Acknowledgments
Financial support from FWO, as well as from the Excellence of Science (EOS) programme (No. 40007501) and the KU Leuven project C14/22/104, is acknowledged. The STFC consolidated grants ST/V001116/1 and ST/P004423/1 and the FNPMLS ERC grant agreement No. 648381 are acknowledged. SGW and RFGR acknowledge funding by the Office of Nuclear Physics, U.S. Department of Energy Grants DE-SC0021176 and DE-SC002117. MAu, MN, AR, JWa, and JWe acknowledge funding from the EU’s H2020-MSCA-ITN Grant No. 861198 ‘LISA’. DH acknowledges financial support from the Swedish Research Council (2020-03505). JL acknowledges financial support from STFC grant ST/V00428X/1. SWB, YCL, WCM, and XFY acknowledge support from the National Natural Science Foundation of China (No. 12350007).

### v2

@@ IV Conclusions
In this work, the measurement of the radiative lifetime of the () state in RaF is presented. The lifetime was measured via delayed multi-step ionization using the CRIS experiment at the CERN-ISOLDE radioactive ion beam facility. The lifetime is determined to be ns. Using this value, the radiative decay rate s-1 is extracted, which further determines the maximum photon scattering rate s-1. The validity of the lifetime measurement in RaF is benchmarked via a measurement of the lifetime of the state in Fr using the same technique. The extracted lifetime for the state in Fr at 83(3) ns in this work is in agreement with the literature value.
The quantities reported in this work are of direct relevance to simulations for efficient laser cooling of RaF. Laser-cooled and trapped neutral molecules are envisioned as sensitives probes for future searches of new physics, and thanks to its exceptionally diagonal Franck-Condon matrix, RaF is expected to scatter 105 photons using only 2 cooling lasers, while using 3 cooling lasers allows scattering an order of magnitude more photons than other laser-cooled molecules. While the photon-scattering acceleration is lower in RaF compared to the lighter homoelectronic molecules, it is shown that the difference becomes insignificant for a molecular source that produces a sufficiently slow beam (50 m/s).
@@ Acknowledgments
Financial support from FWO, as well as from the Excellence of Science (EOS) programme (No. 40007501) and the KU Leuven project C14/22/104, is acknowledged. The STFC consolidated grants ST/V001116/1 and ST/P004423/1 and the FNPMLS ERC grant agreement No. 648381 are acknowledged. SGW and RFGR acknowledge funding by the Office of Nuclear Physics, U.S. Department of Energy Grants DE-SC0021176 and DE-SC002117. MAu, AR, JWa, and JWe acknowledge funding from the EU’s H2020-MSCA-ITN Grant No. 861198 ‘LISA’. DH acknowledges financial support from the Swedish Research Council (2020-03505). JL acknowledges financial support from STFC grant ST/V00428X/1. SWB, YCL, WCM, and XFY acknowledge support from the National Natural Science Foundation of China (No. 12350007).